Framers is in the business of trading in frozen mango pulp tins. It always maintains a



more inventory than required and hence incurs a huge amount of holding (carrying)



costs. It wishes to know the ideal quantity of inventory to be ordered that would



minimize the ordering as well as the holding costs. It provides the following



information:



a. Annual requirement 30,000 packets; cost of one packet is ₹1200; ordering cost is ₹3,240



per order and holding cost is 5 percent. Compute the economic order quantity.



b. What is the total inventory cost if the company has been ordering 2,500 packets with



every order? Would the total inventory cost be higher or lower than the EOQ?

1
Expert's answer
2022-03-18T03:54:02-0400

a) EOQ =square root of: [2(setup costs)(demand rate)] / holding costs.

EOQ = square root of: {[2*3240*30000]/(0.05*1200)}

EOQ = square root of: {[194,400,000/60]}

EOQ = square root of: {3,240,000}

EOQ = 1800 Units.

b) TC = PC + OC + HC,

Where TC is the Total Cost; PC is Purchase Cost; OC is Ordering Cost; and HC is Holding Cost.

At 2500 packets the total cost would be;

TC= 2500(1200)+ 3240 + 0.05(1200*2500)

TC = 3,000,000+3240+150,000

TC = 3,153,240

At 1800 packets, the EOQ, the total cost would be;

TC =1800*1200+3240+0.05*1200*1800

TC = 2,160,000+3240+108,000

TC = 2,271,240

The total inventory cost is higher by ₹882,000 which would be a loss to the company.